# Extended mean field control problems with constraints: The generalized Fritz-John conditions and Lagrangian method

### v1

@@ 3.2 Formulation of C-(E)MFC problems
We make the following assumptions on the model coefficients throughout the section:
@@
    for $i\in I_{1}$ and $j\in I_{2}$, the mappings $b,\sigma,f,g,\phi^{i}$ and $\psi^{j}$ are Borel measurable and continuously differentiable w.r.t. $x\in\mathbb{R}^{n}$. Moreover, $b,\sigma,f,\psi^{j}$ are continuously differentiable w.r.t. $u\in\mathbb{R}^{l}$ and $\phi^{i}$ is continuously Frêchet differentiable w.r.t. $\gamma\in\mathcal{X}$ Moreover, $\phi^{i}$ is jointly continuous in $(t,x)\in[0,T]\times\mathbb{R}^{n}$.
@@
    there is a constant $M>0$ such that, for all $(t,x,u,\gamma)\in[0,T]\times\mathbb{R}^{n}\times\mathbb{R}^{l}\times\mathcal{X}$,
    $\sum_{i\in I_{1}}\left(|\nabla_{x}\phi^{i}(t,x,\gamma)|+\|D_{\gamma}\phi^{i}(t,x,\gamma)\|_{\mathcal{X}}\right)$ $\leq M\left(1+|x|+|u|+\|\gamma\|_{\mathcal{X}}\right),$
    $\sum_{j\in I_{2}}\left(|\nabla_{x}\psi^{j}(t,x,u)|+|\nabla_{u}\psi^{j}(t,x,u)|\right)$ $\leq M.$
@@
    $|\nabla_{x}g(x,\mu)|\leq M(1+|x|+M_{2}(\mu)),$
    with $M_{2}(\mu):=\int_{\mathbb{R}^{n}}|x|^{2}\mu(\mathrm{d}x)<\infty$.
@@
    the mappings $b(t,x,\rho,u),\sigma(t,x,\rho,u),f(t,x,\rho,u)$ and $g(x,\mu)$ are continuously L-differentiable w.r.t. $\rho\in\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$ or $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$. Moreover, there is a constant $M>0$ such that, for all $(t,x,\rho,\mu,u)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}$,
@@
    $\qquad\qquad\leq M(1+|x|+|u|+M_{2}(\rho)+M_{2}(\mu)).$
@@
$\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\rho}b(t,x,\rho,u)(x^{\prime},u^{\prime})|^{2}\rho(\mathrm{d}x^{\prime},\mathrm{d}u^{\prime})\right)^{\frac{1}{2}}$
$\qquad+\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\rho}\sigma(t,x,\rho,u)(x^{\prime},u^{\prime})|^{2}\rho(\mathrm{d}x^{\prime},\mathrm{d}u^{\prime})\right)^{\frac{1}{2}}\leq M,$ (21)
for the positive constant $M$ being given in Assumption 1-(A3).
@@
Compared with Acciaio et al. [1] or Hu et al. [26], we do not need to impose the uniform boundedness on gradients $\nabla_{u}(b,\sigma)$ of coefficients $(b,\sigma)$ of the state dynamics, which is relaxed to be the linear growth condition as in Assumption 1-(A4).
We then consider the following constrained extended MFC problem given by
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[6.00006pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s})\mathrm{d}s$}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s})\mathrm{d}W_{s}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\widetilde{\alpha}^{t}\right)\right]\geq 0$,~{}~{}$i\in I_{1}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\psi^{j}\left(t,X_{t}^{\alpha},\alpha_{t}\right)\geq 0$,~{}~{}$j\in I_{2}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.}\end{cases}$ (22)
@@
Let $i\in I_{1}$ and $(X,\alpha)\in\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}$. Define $g_{i}(t):=\mathbb{E}[\phi^{i}(t,X_{t},\widetilde{\alpha}^{t})]$ with $t\in[0,T]$. Then $g_{i}\in C([0,T];\mathbb{R})$.
@@
$\begin{cases}\displaystyle\mathtt{X}=\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F}}^{2},\\[5.0pt] \displaystyle\mathtt{Y}_{i}=C([0,T];\mathbb{R}),\quad\forall i\in I_{1},\\[5.0pt] \displaystyle\mathtt{Y}_{j}=\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}},\quad\forall j\in I_{2},\\[5.0pt] \displaystyle\mathtt{H}=\mathbb{M}^{2},\end{cases}$ (23)
where, for $x_{1}=(X_{1},\alpha_{1}),x_{2}=(X_{2},\alpha_{2})\in\mathtt{X}$, we define the inner product by
$\langle x_{1},x_{2}\rangle_{\mathtt{X}}=\langle X_{1},X_{2}\rangle_{\mathbb{M}^{2}}+\langle\alpha_{1},\alpha_{2}\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}.$ (24)
Let $\lVert\cdot\rVert_{\mathtt{X}}:=\sqrt{\langle\cdot,\cdot\rangle_{\mathtt{X}}}$ be the induced norm. We also take $\|\cdot\|_{\mathtt{Y}_{i}}$ $(i\in I_{1})$, $\|\cdot\|_{\mathtt{Y}_{j}}$ $(j\in I_{2})$ and $\langle\cdot,\cdot\rangle_{\mathtt{H}}$ as the supremum norm $\|\cdot\|_{\infty}$ on $C([0,T];\mathbb{R}^{n})$, the norm $\|\cdot\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}$ (c.f. (15)) and the inner product $\langle\cdot,\cdot\rangle_{\mathbb{M}^{2}}$ (c.f. (14)), respectively. Furthermore, we define $J:\mathtt{X}\to\mathbb{R}$, $\Phi_{i}:\mathtt{X}\to\mathtt{Y}_{i}$  $(i\in I_{1})$, $\Psi_{j}:\mathtt{X}\to\mathtt{Y}_{j}$ $(j\in I_{2})$ and $F:\mathtt{X}\to\mathtt{H}$ as follows, for $(X,\alpha)\in\mathtt{X}$,
@@
The mappings $\Phi_{i}$ ($i\in I_{1}$) and $\Psi_{j}~{}(j\in I_{2})$ in (25) are well-defined thanks to Assumption 1-(A2) and Lemma 3.5. Thus, we can formulate the following optimization problem with constraints that
@@
are closed and convex cones in $C([0,T];\mathbb{R})$ and $\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}$, respectively. Moreover, one can verify that $K_{1}^{\rm o}=\{f\in C([0,T];\mathbb{R});~{}f(t)>0,~{}\forall t\in[0,T]\}$ and $K_{2}^{\rm o}=\varnothing$. On the other hand, the well-posedness of the controlled Mckean-Vlasov SDE (3.2) implies the following crucial equivalence whose proof is straightforward and hence omitted.
@@ 3.3 FJ optimality condition for abstract problem (26)
We next apply Theorem 2.7 to our C-(E)MFC problem in (26) and derive the FJ optimality condition. To this end, we need to provide the representations of Fréchet differentiablites of the mappings $J,~{}\Phi_{i}~{}(i\in I_{1}),~{}\Psi_{j}~{}(j\in I_{2})$ and $F$ defined in (25). In the sequel, let us set $\rho_{t}=\mathcal{L}(X_{t},\alpha_{t})$ for $t\in[0,T]$.
@@
The mappings $J$, $\Phi_{i}$ ($i\in I_{1}$), $\Psi_{j}~{}(j\in I_{2})$ and $F$ defined in (25) are continuously Frêchet differentiable and their Frêchet derivatives have the following representations that, for $((X,\alpha),L,K)\in\mathtt{X}\times\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}$,
@@
$D_{X}$ $\Phi_{i}(X,\alpha)(L)(t)=-\mathbb{E}[\langle\nabla_{x}\phi^{i}(t,X_{t},\widetilde{\alpha}^{t}),L_{t}\rangle],~{}~{}\forall t\in[0,T],~{}i\in I_{1},$
$D_{X}$ $\Psi_{i}(X,\alpha)(L)(t)=-\int_{0}^{T}\langle\nabla_{x}\psi^{j}(t,X_{t},\alpha_{t}),L_{t}\rangle~{}~{}\forall t\in[0,T],~{}j\in I_{2},$
@@
$D_{\alpha}$ $\Phi_{i}(X,\alpha)(K)(t)=-\mathbb{E}[\langle\nabla_{\beta}\phi^{i}(t,X_{t},\widetilde{\alpha}^{t}),\widetilde{K}^{t}\rangle_{\mathcal{X}}],~{}~{}\forall t\in[0,T],~{}i\in I_{1},$
$D_{\alpha}$ $\Psi_{i}(X,\alpha)(K)(t)=-\langle\nabla_{u}\psi^{j}(t,X_{t},\alpha_{t}),K_{t}\rangle,~{}~{}\forall t\in[0,T],~{}j\in I_{2},$
@@
Then, the following (self-dependent) SDE admits a unique strong solution, for $t\in[0,T]$,
@@
Here, for $t\in[0,T]$, $\overline{Y}_{t}$ stands for the random variable $Y_{t}$ instead of the evaluation $Y_{t}(\omega)$ at $\omega\in\Omega$.
@@
Proof.
To prove the surjection, it suffices to show that, for any $\zeta\in\mathbb{M}^{2}$ admitting the representation $\zeta_{t}=\zeta_{0}+\int_{0}^{t}\zeta^{1}_{s}\mathrm{d}s+\int_{0}^{t}\zeta^{2}_{s}\mathrm{d}W_{s}$ for $t\in[0,T]$, there exists a strong solution $L=(L_{t})_{t\in[0,T]}$ to the SDE, for $t\in[0,T]$,
$L_{t}$ $=\zeta_{0}+\int_{0}^{t}(\nabla_{x}b(s,X_{s},\rho_{s},\alpha_{s})L_{s}+\zeta_{s}^{1})\mathrm{d}s+\int_{0}^{t}(\nabla_{x}\sigma(s,X_{s},\rho_{s},\alpha_{s})L_{s}+\zeta_{s}^{2})\mathrm{d}W_{s}$ (31)
$\quad+\int_{0}^{t}\mathbb{E}^{\prime}\left[\partial_{\mu}b(s,X_{s},\rho_{s},\alpha_{s})(X_{s}^{\prime},\alpha_{s}^{\prime})L_{s}^{\prime}\right]\mathrm{d}s+\int_{0}^{t}\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(s,X_{s},\rho_{s},\alpha_{s})(X_{s}^{\prime},\alpha_{s}^{\prime})L_{s}^{\prime}\right]\mathrm{d}W_{s}.$
In fact, the existence holds if we define $(B,\Sigma)$ in Lemma 3.8 by, for $(t,\omega,x,X)\in[0,T]\times\Omega\times\mathbb{R}^{n}\times L^{2}(\Omega;\mathbb{R}^{n})$,
$\begin{cases}\displaystyle B(t,\omega,x,X):=\nabla_{x}b(t,X_{t}(\omega),\rho_{t},\alpha_{t}(\omega))x+\mathbb{E}^{\prime}\left[\partial_{\mu}b(t,X_{t}(\omega),\rho_{t},\alpha_{t}(\omega))(X_{s}^{\prime},\alpha_{s}^{\prime})X^{\prime}\right]+\zeta_{t}^{1}(\omega),\\[8.00003pt] \Sigma(t,\omega,x,X):=\nabla_{x}\sigma(t,X_{t}(\omega),\rho_{t},\alpha_{t}(\omega))x+\mathbb{E}^{\prime}\left[\partial_{\mu}\sigma(t,X_{t}(\omega),\rho_{t},\alpha_{t}(\omega))(X_{s}^{\prime},\alpha_{s}^{\prime})X^{\prime}\right]+\zeta_{t}^{2}(\omega),\end{cases}$
and by exploiting (3.3), it holds that
$\sup_{(t,x)\in[0,T]\times\mathbb{R}^{n}}\left(\nabla_{x}b(t,X_{t},\rho_{t},\alpha_{t})|+|\nabla_{x}(t,X_{t},\rho_{t},\alpha_{t})|\right)\leq M,~{}{\rm and}$
$\qquad\qquad\qquad\mathbb{E}\left[|\zeta_{0}|^{2}+\int_{0}^{T}\left(|\zeta_{t}^{1}|^{2}+|\zeta_{t}^{2}|^{2}\right)\mathrm{d}t\right]<\infty.$
On the other hand, the injection result holds given that $\ker(D_{X}F(X,\alpha))=0$, which is equivalent to show that the SDE, for $t\in[0,T]$,
$L_{t}=\int_{0}^{t}\nabla_{x}b(s,X_{s},\rho_{s},\alpha_{s})L_{s}\mathrm{d}s+\int_{0}^{t}\nabla_{x}\sigma(s,X_{s},\rho_{s},\alpha_{s})L_{s}\mathrm{d}W_{s}$
admits only a zero solution, which is also a direct consequence of Lemma 3.8. Hence, the proof of the lemma is complete. ∎

### v2

@@ 3.2 Formulation of C-MFC problems
Remark 3.3.
The expectation constraint (13) involving the law $\mathcal{L}(X_{t}^{\alpha})$ is motivated by the variance-type constraint $\mathbb{E}[|X_{t}^{\alpha}+\int_{0}^{t}\alpha_{s}\mathrm{d}s-\mathbb{E}[X_{t}^{\alpha}]|^{2}]\leq a_{t}$, where $t\mapsto a_{t}$ is a deterministic continuous mapping. This type of constraints cannot be encompassed by the pathwise constraint in (14). Moreover, when $\phi^{i}$ is independent of the measure argument $\mu$, the constraint (13) extends the expectation constraint considered by Hu et al. [27] in which the control does not enter into the constraint.
The following assumptions on model coefficients are imposed throughout the section:
@@
    for $i\in I_{1}$ and $j\in I_{2}$, the mappings $b,\sigma,f,g,\phi^{i}$ and $\psi^{j}$ are Borel measurable and continuously differentiable w.r.t. $x\in\mathbb{R}^{n}$. Moreover, $b,\sigma,f,\psi^{j}$ are continuously differentiable w.r.t. $u\in\mathbb{R}^{l}$ and $\phi^{i}$ is continuously Frêchet differentiable w.r.t. $\gamma\in\mathcal{X}$. Moreover, $\phi^{i}$ is jointly continuous in $(t,x,\mu)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})$.
@@
    there is a constant $M>0$ such that, for all $(t,x,\mu,\rho,u,\gamma)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathbb{R}^{l}\times\mathcal{X}$,
    $\sum_{i\in I_{1}}\left(|\nabla_{x}\phi^{i}(t,x,\mu,\gamma)|+\|D_{\gamma}\phi^{i}(t,x,\mu,\gamma)\|_{\mathcal{X}}\right)$ $\leq M\left(1+|x|+M_{2}(\mu)+\|\gamma\|_{\mathcal{X}}\right),$
    $\sum_{j\in I_{2}}\left(|\nabla_{x}\psi^{j}(t,x,\rho,u)|+|\nabla_{u}\psi^{j}(t,x,\rho,u)|\right)$ $\leq M.$
    Here $M_{2}(\mu):=\int_{\mathbb{R}^{n}}|x|^{2}\mu(\mathrm{d}x)<\infty$.
@@
    $|\nabla_{x}g(x,\mu)|\leq M(1+|x|+M_{2}(\mu)).$
@@
    for $i\in I$ and $j\in J$, the mappings $b(t,x,\rho,u),\sigma(t,x,\rho,u),f(t,x,\rho,u)$, $g(x,\mu)$, $\phi^{i}$ and $\psi^{j}$ are continuously L-differentiable w.r.t. $\rho\in\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})$ or $\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})$. Moreover, there is a constant $M>0$ such that, for all $(t,x,\rho,\mu,u,\gamma)\in[0,T]\times\mathbb{R}^{n}\times\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{l})\times\mathcal{P}_{2}(\mathbb{R}^{n})\times\mathbb{R}^{l}\times\mathcal{X}$,
@@
    $~{}+\sum_{i\in I_{1}}\left(\int_{\mathbb{R}^{n}}|\partial_{\mu}\phi^{i}(t,x,\mu,\gamma)(x^{\prime})|^{2}\mu(\mathrm{d}x)\right)^{\frac{1}{2}}\leq M(1+|x|+|u|+M_{2}(\rho)+M_{2}(\mu)+\|\gamma\|_{\mathcal{X}}),$
    and moreover
    $\sum_{j\in I_{2}}\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\rho}\psi^{j}(t,x,\rho,u)(x^{\prime},u^{\prime})|^{2}\rho(\mathrm{d}x^{\prime},\mathrm{d}u^{\prime})\right)^{\frac{1}{2}}\leq M.$
@@
$\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\rho}b(t,x,\rho,u)(x^{\prime},u^{\prime})|^{2}\rho(\mathrm{d}x^{\prime},\mathrm{d}u^{\prime})\right)^{\frac{1}{2}}+\left(\int_{\mathbb{R}^{n}\times\mathbb{R}^{l}}|\partial_{\rho}\sigma(t,x,\rho,u)(x^{\prime},u^{\prime})|^{2}\rho(\mathrm{d}x^{\prime},\mathrm{d}u^{\prime})\right)^{\frac{1}{2}}\leq M,$
for the positive constant $M$ given in Assumption 1-(A3).
@@
Comparing with Acciaio et al. [1] or Hu et al. [27], we do not impose the uniform boundedness on gradients $\nabla_{u}(b,\sigma)$ of coefficients $(b,\sigma)$ of the state dynamics, which is relaxed to be the linear growth condition in Assumption 1-(A4).
We then consider the following C-MFC problem given by
$\begin{cases}\text{minimize $J(\alpha)$ over $\alpha\in\mathscr{U}[0,T]$;}\\[6.00006pt] \displaystyle\text{subject to $X_{t}^{\alpha}=\kappa+\int_{0}^{t}b(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s})\mathrm{d}s$}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$+\int_{0}^{t}\sigma(s,X_{s}^{\alpha},\mathcal{L}(X_{s}^{\alpha},\alpha_{s}),\alpha_{s})\mathrm{d}W_{s}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\mathbb{E}\left[\phi^{i}\left(t,X_{t}^{\alpha},\mathcal{L}(X_{t}^{\alpha}),\widetilde{\alpha}^{t}\right)\right]\geq 0$,~{}~{}$i\in I_{1}$,~{}~{}$\forall t\in[0,T]$,}\\[8.00003pt] \displaystyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{$\psi^{j}\left(t,X_{t}^{\alpha},\mathcal{L}(X_{t}^{\alpha},\alpha_{t}),\alpha_{t}\right)\geq 0$,~{}~{}$j\in I_{2}$,~{}~{}$\mathrm{d}t\times\mathrm{d}\mathbb{P}$-a.s.}\end{cases}$ (16)
@@
Let $i\in I_{1}$ and $(X,\alpha)\in\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}$. Define $g_{i}(t):=\mathbb{E}[\phi^{i}(t,X_{t},\mathcal{L}(X_{t}),\widetilde{\alpha}^{t})]$ with $t\in[0,T]$. Then it holds that $g_{i}\in C([0,T];\mathbb{R})$.
@@
$\mathtt{X}=\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F}}^{2},~{}\mathtt{Y}_{i}=C([0,T];\mathbb{R}),~{}\forall i\in I_{1},~{}\mathtt{Y}_{j}=\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}},~{}\forall j\in I_{2},~{}\mathtt{H}=\mathbb{M}^{2},$ (17)
where, for $x_{1}=(X_{1},\alpha_{1}),x_{2}=(X_{2},\alpha_{2})\in\mathtt{X}$, we define the inner product by $\langle x_{1},x_{2}\rangle_{\mathtt{X}}:=\langle X_{1},X_{2}\rangle_{\mathbb{M}^{2}}+\langle\alpha_{1},\alpha_{2}\rangle_{\mathbb{L}_{\mathbb{F}}^{2}}$. Let $\lVert\cdot\rVert_{\mathtt{X}}:=\sqrt{\langle\cdot,\cdot\rangle_{\mathtt{X}}}$ be the induced norm. We also take $\|\cdot\|_{\mathtt{Y}_{i}}$ $(i\in I_{1})$, $\|\cdot\|_{\mathtt{Y}_{j}}$ $(j\in I_{2})$ and $\langle\cdot,\cdot\rangle_{\mathtt{H}}$ as the supremum norm $\|\cdot\|_{\infty}$ on $C([0,T];\mathbb{R}^{n})$, the norm $\|\cdot\|_{\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}}$ (c.f. (11)) and the inner product $\langle\cdot,\cdot\rangle_{\mathbb{M}^{2}}$ (c.f. (10)), respectively. Furthermore, we define $J:\mathtt{X}\to\mathbb{R}$, $\Phi_{i}:\mathtt{X}\to\mathtt{Y}_{i}$  $(i\in I_{1})$, $\Psi_{j}:\mathtt{X}\to\mathtt{Y}_{j}$ $(j\in I_{2})$ and $F:\mathtt{X}\to\mathtt{H}$ as follows, for $(X,\alpha)\in\mathtt{X}$,
@@
The mappings $\Phi_{i}$ ($i\in I_{1}$) and $\Psi_{j}~{}(j\in I_{2})$ in (18) are well-defined thanks to Assumption 1-(A2) and Lemma 3.6. Thus, we can formulate the optimization problem with constraints that
@@
are closed and convex cones in $C([0,T];\mathbb{R})$ and $\mathbb{L}^{2}_{\mathbb{F},\mathbb{R}}$, respectively. Moreover, one can verify that $K_{1}^{\rm o}=\{f\in C([0,T];\mathbb{R});~{}f(t)>0,~{}\forall t\in[0,T]\}$ and $K_{2}^{\rm o}=\varnothing$. On the other hand, the well-posedness of the controlled Mckean-Vlasov SDE (12) implies the following crucial equivalence whose proof is straightforward and hence omitted.
@@ 3.3 FJ optimality condition for abstract problem (19)
We next apply Theorem 2.6 to our C-MFC problem in (19) and derive the SMP for the necessary optimality condition. To this end, we need to provide the representations of Fréchet differentiablites of the mappings $J,~{}\Phi_{i}~{}(i\in I_{1}),~{}\Psi_{j}~{}(j\in I_{2})$ and $F$ defined in (18). In the sequel, let us set $\mu_{t}=\mathcal{L}(X_{t})$ and $\rho_{t}=\mathcal{L}(X_{t},\alpha_{t})$ for $t\in[0,T]$.
@@
The mappings $J$, $\Phi_{i}$ ($i\in I_{1}$), $\Psi_{j}~{}(j\in I_{2})$ and $F$ defined in (18) are continuously Frêchet differentiable and their Frêchet derivatives have the following representations that, for $(t,(X,\alpha),L,K)\in[0,T]\times\mathtt{X}\times\mathbb{M}^{2}\times\mathbb{L}_{\mathbb{F}}^{2}$, $i\in I_{1}$, and $j\in I_{2}$,
@@
$D_{X}$ $\Phi_{i}(X,\alpha)(L)(t)=-\mathbb{E}\left[\langle\nabla_{x}\phi^{i}(t,X_{t},\mu_{t},\widetilde{\alpha}^{t}),L_{t}\rangle+\mathbb{E}^{\prime}\left[\langle\partial_{\mu}\phi^{i}(t,X_{t},\mu_{t},\widetilde{\alpha}^{t})(X_{t}^{\prime}),L_{t}^{\prime}\rangle\right]\right],$
$D_{X}$ $\Psi_{i}(X,\alpha)(L)(t)=-\langle\nabla_{x}\psi^{j}(t,X_{t},\rho_{t},\alpha_{t}),L_{t}\rangle-\mathbb{E}^{\prime}\left[\langle\partial_{\mu}\psi^{j}(t,X_{t},\rho_{t},\alpha_{t})(X_{t}^{\prime},\alpha_{t}^{\prime}),L_{t}^{\prime}\rangle\right],$
@@
$D_{\alpha}$ $\Phi_{i}(X,\alpha)(K)(t)=-\mathbb{E}\left[\langle D_{\gamma}\phi^{i}(t,X_{t},\mu_{t},\widetilde{\alpha}^{t}),\widetilde{K}^{t}\rangle_{\mathcal{X}}\right],$
$D_{\alpha}$ $\Psi_{i}(X,\alpha)(K)(t)=-\langle\nabla_{u}\psi^{j}(t,X_{t},\rho_{t},\alpha_{t}),K_{t}\rangle-\mathbb{E}^{\prime}\left[\langle\partial_{\nu}\psi^{j}(t,X_{t},\rho_{t},\alpha_{t})(X_{t}^{\prime},\alpha_{t}^{\prime}),K_{t}^{\prime}\rangle\right],$
@@
Then, the following (self-dependent) SDE admits a unique strong solution, for $t\in(0,T]$,
@@
Here, for $t\in[0,T]$, $\overline{Y}_{t}$ stands for the r.v. $Y_{t}$ instead of the evaluation $Y_{t}(\omega)$ at $\omega\in\Omega$.